OpenStudy (anonymous):

Find the zeros for f(x)= (x^2-16x+64)(x-2)

OpenStudy (anonymous):

I can help

OpenStudy (anonymous):

f(x) = x² + 16x + 62 Zeros: f(x) = 0: x² + 16x + 62 = 0 x² + 16x = - 62 x² + 16x + 64 = 64 - 62 (x + 8)² = 2 x + 8 = √2 x = - 8 ± √2 ¯¯¯¯¯¯¯¯¯¯

OpenStudy (anonymous):

Source: 3/11/14

OpenStudy (mathstudent55):

@saynabahmed13 The problem is: (x^2-16x+64)(x-2) = 0 You used the wrong trinomial.

OpenStudy (campbell_st):

you have been given a linear factor and a quadratic factor \[f(x) = (x^2 - 16x + 64)(x - 2)\] the quadratic factor is a perfect square \[x^2 - 2xb + b^2 = (x - b)^2\] can you calculate b?

OpenStudy (campbell_st):

the linear factor (x - 2) to find the zero, let it equal zero and solve for x so x - 2 = 0 what value does x take...?

OpenStudy (campbell_st):

with factoring the perfect square, you need to compare \[x^2 - 16x + 64 ~~~and x^2 -2xb + b^2 \] to find b so it can be written in the factored form \[(x - b)^2\]